# Edge Importance in Complex Networks

### v1

@@ 3.1.1 Network simplification by edge removal
and use $\mathbf{A}-\varepsilon\mathbf{F}$ with $\varepsilon=w_{ij}$.
@@ 3.1.2 Network modification by edge-weight tuning
We describe how to increase the total communicability and use the notation of subsection 3.1.1. The discussion follows [21]. We would like to choose a perturbation $\varepsilon\mathbf{F}$ of $\mathbf{A}$, where $\varepsilon>0$ and $\mathbf{F}$ is of the form (22), so that the Perron root $\rho$ increases as much as possible. This suggests that we choose the indices $i$ and $j$ in (22) so that
@@
We turn to the reduction of the total communicability. Define the matrix $\mathbf{F}$ as above and consider the perturbed matrix $\mathbf{A}-\varepsilon\mathbf{F}$. The parameter $\varepsilon>0$ should be chosen small enough so that this matrix has nonnegative entries only. Moreover, if removing an edge $e_{k}=e(v_{i}\to v_{j})$ associated with of one of the first few of the ordered entries of ${\mathcal{E}}_{\rho}$ results in a disconnected graph and this is undesirable, then we choose $\varepsilon>0$ so that $0<\varepsilon<w_{ij}$. Analogously, if removing the edges $e(v_{i}\leftrightarrow v_{j})$ of an undirected graph ${\mathcal{G}}$ makes the graph disconnected and this is undesirable, then we choose $\varepsilon>0$ so that $\varepsilon<2\,w_{ij}=2\,w_{ji}$.
@@ 3.1.3 Network modification by inclusion of new edges
Let $\mathbf{F}\in{\mathcal{A}}$ be a nonnegative matrix of unit Frobenius norm, $\|\mathbf{F}\|_{F}=1$, and let $\varepsilon>0$ be a small constant. Then
@@
This is the maximal perturbation for the Perron root $\rho$ induced by a unit norm matrix $\mathbf{F}\in{\mathcal{A}}$; see [10, 20]. The quantity
@@
Conversely, if $\kappa_{{\mathcal{A}}}(\rho)\ll\kappa(\rho)$, then the addition of a suitable edge with weight $w_{ij}>0$ (or a suitable pair of edges with weights $w_{ij}=w_{ji}>0$, if the graph is undirected) that increases the ratio $\kappa_{{\mathcal{A}}}(\rho)/\kappa(\rho)$ may be an appropriate strategy to increase the Perron communicability. Recall that $\widehat{\mathcal{A}}$ denotes the cone of the nonnegative matrices in $\mathbb{R}^{n\times n}$ whose sparsity structure is given by the zero entries of $\mathbf{A}$ except for the diagonal entries. Perturb the entry $w_{ij}$ with $i\neq j$ of $\mathbf{A}$ and let
@@
for some index pair $\{i,j\}$. The entries of Wilkinson perturbation $\mathbf{W}$ reveal which edges should be added to the network graph to increase the communicability, namely edges whose associated entries of the matrix $\mathbf{W}$ are large. The procedure for the construction of the vector ^$\widehat{{\mathcal{E}}}_{\rho}\in\mathbb{R}^{n^{2}-n-m}$ that gives the Perron virtual importance of the virtual edges is the following:
@@
3.2 Network modification criteria for large-sized networks
@@
If our aim is to perturb or set to zero suitable positive entries of $\mathbf{A}\in{\mathcal{A}}$, then the Wilkinson perturbation $\mathbf{W}$ does not have to be constructed, since one only needs the entries of $\mathbf{W}|_{\mathcal{A}}\in{\mathcal{A}}$. The Perron edge importance vector ${\mathcal{E}}_{\rho}\in\mathbb{R}^{m}$, relevant to the $m$ edges of ${\mathcal{G}}$, can be evaluated as discussed in Subsection 3.1.1.
@@ 4.1 Medium-sized networks
EXAMPLE 4.1.
Consider the adjacency matrix $\mathbf{A}\in\mathbb{R}^{500\times 500}$ for the network Air500 in [9]. This data set describes flight connections for the top 500 airports worldwide based on total passenger volume. The flight connections between airports are for the year from 1 July 2007 to 30 June 2008. The network is represented by a directed unweighted connected graph ${\mathcal{G}}$ with $n=500$ vertices and $m=24009$ directed edges. The vertices of the network are the airports and the edges represent direct flight routes between two airports.
@@
As for both the edge importance vector ${\mathcal{E}}_{L_{f}}$ and the virtual edge importance vector ^$\widehat{\mathcal{E}}_{L_{f}}$, the same results displayed below are obtained regardless of whether $\nabla T_{\mathcal{G}}$ or $\widetilde{\nabla T}_{\mathcal{G}}$ is used.
The Perron communicability in (20) is $P_{\mathcal{G}}=1.9132\cdot 10^{38}$. The Perron root and left and right Perron vectors have been evaluated by using the MATLAB function eig.
@@
Table 1: Example 4.1. The first $10$ flight connections to remove, in order to reduce the complexity of Air500 without changing the network communication significantly, according to determination of the edge importance based on gradient and on Perron root sensitivity, respectively.
In Table 1 the $10$ smallest entries of both the edge importance vector ${\mathcal{E}}_{L_{f}}$ and the Perron edge importance vector ${\mathcal{E}}_{\rho}$ are shown, along with the corresponding edges. This is useful for determining which edges to remove in order to reduce the complexity of the Air500 network (cf. Subsections 2.1.1 and 3.1.1). One can observe that the elimination of the air connection from Santos Dumont Airport - Rio de Janeiro, Brazil (SDU) to Congonhas Airport - S. Paulo, Brazil (CGH), at the third position in the ranking given by ${\mathcal{E}}_{\rho}$, would disconnect the network. We therefore only remove the two edges in bold face in Table 1, i.e., the two most irrelevant edges - according to edge importance determination based on both gradient and Perron root sensitivity - which correspond to the flight connection between San Antonio International Airport - San Antonio, Texas (TSA) and Penghu Airport, Taiwan (MZG). This results in the network ${\mathcal{G}}_{1}$, for which we have:
@@
Table 2: Example 4.1. The first $10$ flight connections to increase/decrease, in order to increase/decrease the network communication in Air500, according to determination of the edge importance based on gradient and on Perron root sensitivity, respectively.
Table 2 shows the $10$ largest entries of both the edge importance vector ${\mathcal{E}}_{L_{f}}$ and the Perron edge importance vector ${\mathcal{E}}_{\rho}$ along with the corresponding edges. In order to obtain a relatively large reduction in the total communicability, we set to zero the weights associated with the two largest entries of both ${\mathcal{E}}_{L_{f}}$ and ${\mathcal{E}}_{\rho}$. This means we remove the edges that represent air route between John F. Kennedy International Airport - New York City (JFK) and Atlanta Hartsfield-Jackson Airport, Georgia (ATL). This results in the network ${\mathcal{G}}_{2}$, for which, as it is apparent, the reduction in the total communicability is much larger than in ${\mathcal{G}}_{1}$:
@@
Following the discussion in Subsections 2.1.2 and 3.1.2, in order to obtain a relatively large increase in the total communicability, we increase by $1$ the edge-weights associated with the two largest entries of both ${\mathcal{E}}_{L_{f}}$ and ${\mathcal{E}}_{\rho}$ (i.e., the air route between John F. Kennedy International Airport - New York City (JFK) and Atlanta Hartsfield-Jackson Airport, Georgia (ATL)). For the so obtained network ${\mathcal{G}}_{3}$, one has
@@
Table 3: Example 4.1. The first $10$ flight connections that should be added, in order to enhance the network communication in Air500, according to determination of the virtual edge importance based on gradient and on Perron root sensitivity, respectively.
Finally, following the discussion in Subsections 2.1.3 and 3.1.3, we display in Table 3 the $10$ largest entries of the total virtual edge importance vector ^$\widehat{\mathcal{E}}_{L_{f}}$ and the $10$ largest entries of the Perron virtual edge importance vector ^$\widehat{{\mathcal{E}}}_{\rho}$, along with the corresponding nonexistent edges. Notice that the edges associated with the two largest entries of both ^$\widehat{\mathcal{E}}_{L_{f}}$ and ^$\widehat{{\mathcal{E}}}_{\rho}$ cannot be considered because they consist in the missing air route between John F. Kennedy International Airport - New York City (JFK) and LaGuardia Airport - New York City (LGA). While taking note of the suggestion to strengthen or create a shuttle service bus between such airports, we proceed to consider the third and the fourth best nonexistent edges according to determination of the edge importance based on gradient, that is to say the routes from Heathrow Airport - London, England (LHR) to Atlanta Hartsfield-Jackson Airport, Georgia (ATL) and from Amsterdam Schiphol Airport, Netherlands (AMS) to Dallas/Fort Worth International Airport, Texas (DFW), and set their weights to $1$. This way, we obtain the network ${\mathcal{G}}_{4}$, for which one has:
@@
Conversely, setting to $1$ the weight of the third and the fourth best nonexistent edges according to ^$\widehat{{\mathcal{E}}}_{\rho}$, i.e., from Midway International Airport - Chicago, Illinois (MDW) to John F. Kennedy International Airport - New York City (JFK) and from Heathrow Airport - London, England (LHR) to Hartsfield-Jackson Airport - Atlanta, Georgia (ATL), we get the network ${\mathcal{G}}_{5}$, where:
@@
EXAMPLE 4.2.
Consider the undirected unweighted graph $\mathcal{G}$ that represents the German highway system network Autobahn. The graph, which is available at [9], has $n=1168$ vertices representing German locations and $m=1243$ edges representing highway segments that connect them. Therefore, the adjacency matrix $\mathbf{A}\in\mathbb{R}^{1168\times 1168}$ for this network has $2486$ nonvanishing entries.
@@
The slight difference in the edge importance vectors computed using (9) or (10) gives rise to a different ordering of the edges corresponding with their $10$ smallest entries (which in fact differ of ${\cal O}(10^{-12})$) as displayed in Table 4. Despite the fact that removing the two edges in bold face in the second column of Table 4 results in the (disconnected) network ${\mathcal{G}}_{1}$ (see Figure 1(a)) while removing the two edges in bold face in the third column of Table 4 returns the disconnected network ${\widetilde{\mathcal{G}}}_{1}$ (see Figure 1(b)), however we have the same results for both ${\mathcal{G}}_{1}$ and ${\widetilde{\mathcal{G}}}_{1}$:
@@
h!tb Refer to caption Refer to caption (a) (b) Refer to caption (c)
Figure 1: Example 4.2. The vertices (marked in red) that are connected by the edges to be removed in order to simplify the network according to ${\mathcal{E}}_{L_{f}}$ for (a), according to ~${\widetilde{\mathcal{E}}}_{L_{f}}$ for (b), and according to ${\mathcal{E}}_{\rho}$ for (c).
Conversely, the largest entries of the edge importance vectors and the virtual edge importance vectors computed using (9) or (10) correspond to the same edges displayed in Table 5. Setting to zero the weights associated with the two highway segments Duisburg - Düsseldorf and München - Kirchheim (both in bold face in the first column of Table 5) results in the network ${\mathcal{G}}_{2}$, for which one has
@@
On the other hand, setting to one the (vanishing) entries of $A$ associated with the two virtual highway segments München - Duisburg and München - Hamburg (both in bold face in the second column of Table 5) returns the network ${\mathcal{G}}_{4}$ with
$T_{{\mathcal{G}}_{4}}=1.3662\cdot 10^{4};\;\;\;\;\frac{T_{{\mathcal{G}}_{4}}-T_{\mathcal{G}}}{T_{\mathcal{G}}}=8.7477\cdot 10^{-2}.\;\;\;\;\;\;\;\;$
As for Perron communicability in Autobahn, one has $P_{\mathcal{G}}=2.2448\cdot 10^{3}$. Although the graph ${\mathcal{G}}$ is irreducible, some entries of the Perron vector $x$ are close to machine precision. In particular, the edges to remove in order to simplify the network, that are associated with the two smallest entries of the Perron edge importance vector, are the final highway segments Wildsdruff - Wildeck and Wüstenbrand - Wommen, connecting the four vertices associated with such quasi-zero components of $x$ (see Figure 1(c)). This results in the (disconnected) network ${\mathcal{G}}_{5}$, for which we have
$P_{{\mathcal{G}}_{5}}=2.2448\cdot 10^{3};\;\;\;\;\frac{P_{{\mathcal{G}}}-P_{{\mathcal{G}_{5}}}}{P_{\mathcal{G}}}=-4.8619\cdot 10^{-15};\;\;\;\;\;\;\;\;T_{{\mathcal{G}}_{5}}=1.2547\cdot 10^{4};\;\;\;\;\frac{T_{{\mathcal{G}}}-T_{{\mathcal{G}}_{5}}}{T_{\mathcal{G}}}=1.2521\cdot 10^{-3}.$
So the simplification is slightly less satisfactory than that obtained in ${\mathcal{G}}_{1}$ or in ${\widetilde{\mathcal{G}}}_{1}$. Elimination of the edges associated with the two highway segments Duisburg - Düsseldorf and Essen - Duisburg (both in bold face in the first column of Table 6) returns the network ${\mathcal{G}}_{6}$, for which we
$P_{{\mathcal{G}}_{6}}=1.2677\cdot 10^{3};\;\;\;\;\frac{P_{{\mathcal{G}}}-P_{{\mathcal{G}}_{6}}}{P_{\mathcal{G}}}=4.3527\cdot 10^{-1};\;\;\;\;\;\;\;\;T_{{\mathcal{G}}_{6}}=1.2155\cdot 10^{4};\;\;\;\;\frac{T_{{\mathcal{G}}}-T_{{\mathcal{G}}_{6}}}{T_{\mathcal{G}}}=3.2513\cdot 10^{-2}.$
Again the decrease is less than that in ${\mathcal{G}}_{2}$. Conversely, increasing by one the weights associated with the same edges results in the network ${\mathcal{G}}_{7}$, for which one has
$P_{{\mathcal{G}}_{7}}=2.5248\cdot 10^{3};\;\;\;\;\frac{P_{{\mathcal{G}}_{7}}-P_{{\mathcal{G}}}}{P_{\mathcal{G}}}=1.2472\cdot 10^{-2};\;\;\;\;\;\;\;\;T_{{\mathcal{G}}_{7}}=1.3480\cdot 10^{4};\;\;\;\;\frac{T_{{\mathcal{G}}_{7}}-T_{{\mathcal{G}}}}{T_{\mathcal{G}}}=7.2982\cdot 10^{-2}.$
The increase in this case is greater than that in ${\mathcal{G}}_{3}$. Finally, setting to one the (vanishing) entries of $A$ associated with the two virtual highway segments Essen - Düsseldorf and Gelsenkirchen - Duisburg (both in bold face in the second column of Table 6) results in the network ${\mathcal{G}}_{8}$ with
$P_{{\mathcal{G}}_{8}}=2.4892\cdot 10^{3};\;\;\;\;\frac{P_{{\mathcal{G}}_{8}}-P_{{\mathcal{G}}}}{P_{\mathcal{G}}}=1.0885\cdot 10^{-1};\;\;\;\;\;\;\;\;T_{{\mathcal{G}}_{8}}=1.3271\cdot 10^{4};\;\;\;\;\frac{T_{{\mathcal{G}}_{8}}-T_{{\mathcal{G}}}}{T_{\mathcal{G}}}=5.6327\cdot 10^{-2}.$
Therefore the result is less satisfactory than that obtained for the network ${\mathcal{G}}_{4}$.
Notice that one has $\|\mathbf{W}|_{{\mathcal{A}}}\|_{F}=\sqrt{2}\|\mathbf{W}|_{{\mathcal{L}}}\|_{F}=5.1765\cdot 10^{-1}$.
@@
Table 4: Example 4.2. The first $10$ highway segments that could be removed, in order to reduce the complexity of Autobahn without changing the network communication significantly, according to determination of the edge importance based on gradient. The edges in the second and in the third column are determined using (9) and (10), respectively.
${\mathcal{E}}_{L_{f}}$ $e(v_{i}\leftrightarrow v_{j})$${\widehat{\mathcal{E}}}_{L_{f}}$ $e(v_{i}\leftrightarrow v_{j})$

### v2

@@ 3.1.1 Network simplification by edge removal
and use $\mathbf{A}-\varepsilon\mathbf{F}$ with $\varepsilon=w_{ij}$ and $\mathbf{F}$ given by (23),
@@ 3.1.2 Network modification by edge-weight tuning
We describe how to increase the total communicability and use the notation of Subsection 3.1.1. The discussion follows NR . We would like to choose a perturbation $\varepsilon\mathbf{F}$ of $\mathbf{A}$, where $\varepsilon>0$ and $\mathbf{F}$ is of the form (23), so that the Perron root $\rho$ increases as much as possible. This suggests that we choose the indices $i$ and $j$ in (23) so that
@@
We turn to the reduction of the total communicability. Define the matrix $\mathbf{F}$ as above and consider the perturbed matrix $\mathbf{A}-\varepsilon\mathbf{F}$. The parameter $\varepsilon>0$ should be chosen small enough so that this matrix has nonnegative entries only. Moreover, if removing an edge $e_{k}=e(v_{i}\to v_{j})$ associated with one of the first few of the ordered entries of ${\mathcal{E}}_{\rho}$ results in a disconnected graph and this is undesirable, then we choose $\varepsilon>0$ so that $0<\varepsilon<w_{ij}$. Analogously, if removing the edges $e(v_{i}\leftrightarrow v_{j})$ of an undirected graph ${\mathcal{G}}$ makes the graph disconnected and this is undesirable, then we choose $\varepsilon>0$ so that $\varepsilon<2\,w_{ij}=2\,w_{ji}$.
@@ 3.1.3 Network modification by inclusion of new edges
Let $\mathbf{F}\in{\mathcal{A}}$ be a nonnegative matrix with $\|\mathbf{F}\|_{F}=1$, and let $\varepsilon>0$ be a small constant. Then
@@
This is the maximal perturbation for the Perron root $\rho$ induced by a unit norm matrix $\mathbf{F}\in{\mathcal{A}}$; see EHNR ; NP . The quantity
@@
Conversely, if $\kappa_{{\mathcal{A}}}(\rho)\ll\kappa(\rho)$, then the addition of a suitable edge with weight $w_{ij}>0$ (or a suitable pair of edges with weights $w_{ij}=w_{ji}>0$, if the graph is undirected) that increases the ratio $\kappa_{{\mathcal{A}}}(\rho)/\kappa(\rho)$ may be an appropriate strategy to increase the Perron communicability. Recall that $\widehat{\mathcal{A}}$ denotes the cone of the nonnegative matrices in $\mathbb{R}^{n\times n}$ whose sparsity structure is given by the zero entries of $\mathbf{A}$ except for the diagonal entries. Perturb the entry $w_{ij}$ with $i\neq j$ of $\mathbf{A}$ and assume that
@@
for some index pair $\{i,j\}$. The entries of the Wilkinson perturbation $\mathbf{W}$ reveal which edges should be added to the network to increase the communicability, namely edges whose associated entries of the matrix $\mathbf{W}$ are large. The procedure for the construction of the vector ^$\widehat{{\mathcal{E}}}_{\rho}\in\mathbb{R}^{n^{2}-n-m}$ that gives the Perron virtual importance of the virtual edges is the following:
@@
3.2 Network modification criteria for large-scale networks
@@
If our aim is to perturb or set to zero suitable positive entries of $\mathbf{A}\in{\mathcal{A}}$, then the Wilkinson perturbation $\mathbf{W}$ does not have to be constructed, since one only needs the entries of $\mathbf{W}|_{\mathcal{A}}\in{\mathcal{A}}$. The Perron edge importance vector ${\mathcal{E}}_{\rho}\in\mathbb{R}^{m}$, associated with the $m$ edges of ${\mathcal{G}}$, can be evaluated as discussed in Subsection 3.1.1.
@@ 4 Computed examples
4.1 A synthetic example
The following example explains why we might be interested in estimating the total communicability without evaluating the exponential of the adjacency matrix associated with the given network.
Example 1
Consider the undirected and connected graph where each vertex represents a person in a line. Each person can communicate only with the following and the preceding persons. The adjacency matrix associated with such a network is the symmetric tridiagonal Toeplitz matrix
$\mathbf{A}=\left[\begin{array}[]{cccccccc}\phantom{vv}&\sigma&&&&&&\\ \sigma&\phantom{vv}&\sigma&&&&&\\ &\sigma&\phantom{vv}&\cdot&&&&\\ &&\cdot&\phantom{vv}&\cdot&&\\ &&&\cdot&\phantom{vv}&\sigma\\ &&&&\sigma&\phantom{vv}\end{array}\right]\in\mathbb{R}^{n\times n},$ (26)
with $\sigma>0$. It is intuitive that the closer the vertices are to the center of the graph, the better connected they are and, therefore, the more important. Also, it is immediate to observe that the best strategy to improve communication in this network is to make the two vertices at the ends adjacent by adding the new undirected edge $e(v_{1}\leftrightarrow v_{n})$ with weight $\sigma$, which makes the adjacency matrix a circulant. This scenario is consistent with the structure of the Wilkinson matrix $\mathbf{W}$ associated with the Perron root of $\mathbf{A}$, which we show in Figure 1(b). This matrix is independent of $\sigma$; see (NP, , Proposition 1). Therefore the network modifications based on the Perron root sensitivity described in Section 3 are in agreement with the previous observations and considerations.
Instead, due to round-off errors introduced in the representation and evaluation of the matrix exponential, we have that the the structure of the matrix $L_{\exp_{0}}(\mathbf{A}^{T},ee^{T})$ (see Figure 1(a)) does not completely reflect the scenario described above. In fact, the techniques based on the gradient (3) described in Section 2 do not identify the same edge as the Wilkinson perturbation. For instance, let $n$ be even and $\sigma=1$. Then we obtain for $n\geq 36$ that the computed gradient does not identify the undirected edge $e(v_{n/2}\leftrightarrow v_{n/2+1})$ as the most important one.
Refer to caption Refer to caption
(a) (b)
Figure 1: Example 1. Structure of $L_{\exp_{0}}(\mathbf{A},ee^{T})$ (left picture (a)) and of $\mathbf{W}$ (right picture (b)) for the matrix $\mathbf{A}\in\mathbb{R}^{50\times 50}$ in (26) with $\sigma=1$.
@@ 4.2 Medium-sized networks
Example 2
Consider the adjacency matrix $\mathbf{A}\in\mathbb{R}^{500\times 500}$ for the network Air500 in air500_autobahn . This data set describes flight connections for the top 500 airports worldwide based on total passenger volume. The flight connections between airports are for the year from 1 July 2007 to 30 June 2008, and the network is represented by a directed unweighted connected graph ${\mathcal{G}}$ with $n=500$ vertices and $m=24009$ directed edges. Vertices of the network are airports and edges represent direct flight routes between two airports.
@@
For both the edge importance vector ${\mathcal{E}}_{L_{f}}$ and the virtual edge importance vector ^$\widehat{\mathcal{E}}_{L_{f}}$, the same results are obtained regardless of whether $\nabla T_{\mathcal{G}}$ or $\widetilde{\nabla T}_{\mathcal{G}}$ is used; see below. We remark that evaluating $L_{f}(\mathbf{A}^{T},ee^{T})$ by (9) required the average time $t_{1}\approx 3.4\cdot 10^{-1}$ in $10^{4}$ tests, while evaluating $L_{f}(\mathbf{A}^{T},ee^{T})$ using (10) required the average time $t_{2}\approx 1.5\cdot 10^{-1}$ in the same tests, with a relative average time saving
$\frac{t_{1}-t_{2}}{t_{1}}\approx 5.6\cdot 10^{-1}.$
The Perron communicability in (21) is $P_{\mathcal{G}}=1.9132\cdot 10^{38}$. The Perron root and left and right Perron vectors were evaluated with the MATLAB function eig.
@@
Table 1: Example 2. The first $10$ flight connections to remove in order to reduce the complexity of Air500 without changing the network communicability significantly, according to the determination of the edge importance based on gradient and on Perron root sensitivity, respectively.
In Table 1 the $10$ smallest entries of both the edge importance vector ${\mathcal{E}}_{L_{f}}$ and the Perron edge importance vector ${\mathcal{E}}_{\rho}$ are shown along with the corresponding edges. This is useful for determining which edges to remove in order to reduce the complexity of the Air500 network (cf. Subsections 2.1.1 and 3.1.1). One can observe that the elimination of the connection from Santos Dumont Airport, Rio de Janeiro, Brazil (SDU) to Congonhas Airport, S. Paulo, Brazil (CGH), at the third position in the ranking given by ${\mathcal{E}}_{\rho}$, would disconnect the network. We therefore only remove the two edges in bold face in Table 1, i.e., the two most irrelevant edges - according to edge importance determination based on both gradient and Perron root sensitivity. These edges correspond to the flight connection between San Antonio International Airport, San Antonio, Texas (TSA) and Penghu Airport, Taiwan (MZG). This results in the network ${\mathcal{G}}_{1}$, for which we have:
@@
Table 2: Example 2. The first $10$ flight connections to increase/decrease in order to increase/decrease the network communicability in Air500 according to the determination of the edge importance based on gradient and on Perron root sensitivity, respectively.
Table 2 shows the $10$ largest entries of both the edge importance vector ${\mathcal{E}}_{L_{f}}$ and the Perron edge importance vector ${\mathcal{E}}_{\rho}$ along with the corresponding edges. In order to obtain a relatively large reduction in the total communicability, we set to zero the weights associated with the two largest entries of both ${\mathcal{E}}_{L_{f}}$ and ${\mathcal{E}}_{\rho}$. This means we remove the edges that represent air route between John F. Kennedy International Airport, New York City (JFK) and Atlanta Hartsfield-Jackson Airport, Georgia (ATL). This results in the network ${\mathcal{G}}_{2}$, for which, as it is apparent, the reduction in the total communicability is much larger than in ${\mathcal{G}}_{1}$:
@@
Following the discussion in Subsections 2.1.2 and 3.1.2, in order to obtain a relatively large increase in the total communicability, we increase by $1$ the edge-weights associated with the two largest entries of both ${\mathcal{E}}_{L_{f}}$ and ${\mathcal{E}}_{\rho}$ (i.e., the air route between John F. Kennedy International Airport, New York City (JFK), and Atlanta Hartsfield-Jackson Airport, Georgia (ATL)). For the so obtained network ${\mathcal{G}}_{3}$, one has
@@
Table 3: Example 2. The first $10$ flight connections that should be added, in order to enhance the communicability in Air500, according to determination of the virtual edge importance based on gradient and on Perron root sensitivity, respectively.
Finally, following the discussion in Subsections 2.1.3 and 3.1.3, we display in Table 3 the $10$ largest entries of the total virtual edge importance vector ^$\widehat{\mathcal{E}}_{L_{f}}$ and the $10$ largest entries of the Perron virtual edge importance vector ^$\widehat{{\mathcal{E}}}_{\rho}$, along with the corresponding nonexistent edges. Notice that the edges associated with the two largest entries of both ^$\widehat{\mathcal{E}}_{L_{f}}$ and ^$\widehat{{\mathcal{E}}}_{\rho}$ cannot be considered because they model the missing air route between John F. Kennedy International Airport, New York City, (JFK), and La Guardia Airport, New York City, (LGA), and are too close to justify a flight route. The entries of the table suggest that there should be a shuttle service between these vertices and, indeed, such a shuttle service exists. We proceed to consider the third and fourth best nonexistent edges according to the edge importance based on the gradient, that is we consider the routes from Heathrow Airport, London, England, (LHR) to Atlanta Hartsfield-Jackson Airport, Georgia, (ATL) and from the Amsterdam Schiphol Airport, Netherlands, (AMS) to Dallas/Fort Worth International Airport, Texas, (DFW), and set their weights to $1$. This way, we obtain the network ${\mathcal{G}}_{4}$, for which one has
@@
Conversely, setting to $1$ the weight of the third and the fourth best nonexistent edges according to ^$\widehat{{\mathcal{E}}}_{\rho}$, i.e., the edges that represent flights from Midway International Airport, Chicago, Illinois, (MDW) to John F. Kennedy International Airport, New York City, (JFK) and from Heathrow Airport, London, England, (LHR) to Hartsfield-Jackson Airport, Atlanta, Georgia (ATL), we obtain the network ${\mathcal{G}}_{5}$ with
@@
Example 3
Consider the undirected unweighted graph $\mathcal{G}$ that represents the German highway system network Autobahn. The graph, which is available at air500_autobahn , has $n=1168$ vertices representing German locations and $m=1243$ edges that represent highway segments connecting them. Therefore, the adjacency matrix $\mathbf{A}\in\mathbb{R}^{1168\times 1168}$ for this network has $2486$ nonvanishing entries.
@@
Notice that evaluating the matrix $L_{f}(\mathbf{A}^{T},ee^{T})$ in $10^{4}$ tests required the average time $t_{1}\approx 1.8\cdot 10^{0}$ using (9) and the average time $t_{2}\approx 3.2\cdot 10^{-1}$ using (10), with a relative average time saving
$\frac{t_{1}-t_{2}}{t_{1}}\approx 8.2\cdot 10^{-1}.$
The slight difference in the edge importance vectors computed using (9) or (10) gives rise to different orderings of the edges corresponding to the $10$ smallest entries (which in fact differ about ${\cal O}(10^{-12})$). This is displayed in Table 4. Removing the two edges in bold face in the second column of Table 4 results in the (disconnected) network ${\mathcal{G}}_{1}$ (see Figure 2(a)), while removing the two edges in bold face in the third column of Table 4 yields the disconnected network ${\widetilde{\mathcal{G}}}_{1}$ (see Figure 2(b)). Nevertheless, we have the same result for both the graphs ${\mathcal{G}}_{1}$ and ${\widetilde{\mathcal{G}}}_{1}$:
@@
Refer to caption Refer to caption
(a) (b)
Figure 2: Example 3. The vertices (marked in red) that are connected by the edges to be removed in order to simplify the network according to ${\mathcal{E}}_{L_{f}}$ for (a), according to ~${\widetilde{\mathcal{E}}}_{L_{f}}$ for (b), and according to ${\mathcal{E}}_{\rho}$ for (c).
Conversely, the largest entries of the edge importance vectors computed using (9) or (10) correspond to the same edges displayed in the first column of Table 5. Setting to zero the weights associated with the two highway segments Duisburg - Düsseldorf and München - Kirchheim (both in bold face in the first column of Table 5) results in the network ${\mathcal{G}}_{2}$, for which one has
@@
As for Perron communicability in Autobahn, one has $P_{\mathcal{G}}=2.2448\cdot 10^{3}$. Although the graph ${\mathcal{G}}$ is irreducible, some entries of the Perron vector $x$ are close to machine precision. In particular, the edges to remove in order to simplify the network, that are associated with the two smallest entries of the Perron edge importance vector, are the highway segments Wildsdruff - Wildeck and Wüstenbrand - Wommen, which connect four vertices associated with such quasi-zero components of $x$ (see Figure 2(c)). This results in the (disconnected) network ${\mathcal{G}}_{4}$, for which we have
$P_{{\mathcal{G}}_{4}}=2.2448\cdot 10^{3};\;\;\;\;\frac{P_{{\mathcal{G}}}-P_{{\mathcal{G}_{4}}}}{P_{\mathcal{G}}}=-4.8619\cdot 10^{-15};\;\;\;\;\;\;\;\;T_{{\mathcal{G}}_{4}}=1.2547\cdot 10^{4};\;\;\;\;\frac{T_{{\mathcal{G}}}-T_{{\mathcal{G}}_{4}}}{T_{\mathcal{G}}}=1.2521\cdot 10^{-3}.$
Hence, this simplification is slightly less satisfactory than the ones described by ${\mathcal{G}}_{1}$ and ${\widetilde{\mathcal{G}}}_{1}$. Elimination of the edges associated with the two highway segments Duisburg - Düsseldorf and Essen - Duisburg (both in bold face in the second column of Table 5) gives the network ${\mathcal{G}}_{5}$, for which we have
$P_{{\mathcal{G}}_{5}}=1.2677\cdot 10^{3};\;\;\;\;\frac{P_{{\mathcal{G}}}-P_{{\mathcal{G}}_{5}}}{P_{\mathcal{G}}}=4.3527\cdot 10^{-1};\;\;\;\;\;\;\;\;T_{{\mathcal{G}}_{5}}=1.2155\cdot 10^{4};\;\;\;\;\frac{T_{{\mathcal{G}}}-T_{{\mathcal{G}}_{5}}}{T_{\mathcal{G}}}=3.2513\cdot 10^{-2}.$
Again the decrease is less than for ${\mathcal{G}}_{2}$. Conversely, increasing by one the weights associated with the same edges results in the network ${\mathcal{G}}_{6}$, for which one has
$P_{{\mathcal{G}}_{6}}=2.5248\cdot 10^{3};\;\;\;\;\frac{P_{{\mathcal{G}}_{6}}-P_{{\mathcal{G}}}}{P_{\mathcal{G}}}=1.2472\cdot 10^{-2};\;\;\;\;\;\;\;\;T_{{\mathcal{G}}_{6}}=1.3480\cdot 10^{4};\;\;\;\;\frac{T_{{\mathcal{G}}_{6}}-T_{{\mathcal{G}}}}{T_{\mathcal{G}}}=7.2982\cdot 10^{-2}.$
The increase in this case is greater than for ${\mathcal{G}}_{3}$.
We do not address the issue of adding new highway segments, because the feasibility of building new highway segments depends on many issues that are not included in our model such as the length of the new segments and properties of the regions, e.g., the presence of mountains, valleys and lakes, that the new segments would traverse.
@@
Table 4: Example 3. The first $10$ highway segments that could be removed in order to reduce the complexity of Autobahn without changing the network communicability significantly, according to the determination of the edge importance based on the gradient. The edges in the second and third columns are determined by (9) and (10), respectively.
${\mathcal{E}}_{L_{f}}$ $e(v_{i}\leftrightarrow v_{j})$ ${\mathcal{E}}_{\rho}$ $e(v_{i}\leftrightarrow v_{j})$